Thirty students bought pennants for the football game. Plain pennants cost $4 each and fancy ones cost $8 each. If the total bill was $168, how many students bought the fancy pennants?
Answer by checkley77(12844)

P+F=30
P=30-F
4P+8F=168
4(30-F)+8F=168
120-4F+8F=168
4F=168-120
4F=48
F=48/4
F=12 FANCY PENNANTS WERE SOLD.
30-12=18 PLAIN PENNANTS WERE SOLD.
PROOF:
4*18+8*12=168
72+96=168
168=168